OpenStudy (anonymous):

help with dividing binomials please ( t^3 - t^2 + t + 1) / (t + 1)

OpenStudy (anonymous):

I'd go synthetic on this one.

OpenStudy (anonymous):



OpenStudy (anonymous):

All your coefficients are either +1 or -1 so polynomial long division would easy enough too, but synthetic division is so much cleaner.

OpenStudy (anonymous):

so final answer is t^2-2t+3-(2/t^3-t^2+t+1) right?

OpenStudy (anonymous):

i do have t^2 - 2t +3 but i wasnt sure how to get the rest

OpenStudy (anonymous):

then since the remainder isnt 0 u hav to divide the remainder by the question so its t^2-2t+3-(2/t^3-t^2+t+1) right? understand?

OpenStudy (anonymous):

the parentheses is considered the remainder?

OpenStudy (anonymous):

if the remainder was 0 u would leave it like wat u said

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ahhh i think i see

OpenStudy (anonymous):

but u can write it as a fraction with no paranthesis the remainder is 2 with the synthetic division but then we divide by thwe question so its like i wrote

OpenStudy (anonymous):

thanks i get it

OpenStudy (anonymous):

no problemo

OpenStudy (anonymous):

Er, there's something wrong there. You express the remainder as -2 over the divisor, not over the dividend...

OpenStudy (anonymous):

( t^3 - t^2 + t + 1) / (t + 1)

OpenStudy (anonymous):

Summary \[ \frac{t^3-t^2+t+1}{t+1}=t^2-2t+3 -\frac{2}{t+1} \]

OpenStudy (anonymous):

@CliffSedge's amswer is correct.\[t^2-2 t+3-\frac{2}{t+1} \]

OpenStudy (anonymous):

yes